OpenStudy (anonymous):

simplify the expression 15 3/4 = t + 4 5/8 will give medal to best answer.

OpenStudy (anonymous):

\[15 \frac{ 3 }{ 4 } = t + 4 \frac{ 5 }{ 8 }\]

OpenStudy (anonymous):

15 3/4 = t + 4 5/8 15 6/8 = t + 4 5/8 - 4 5/8 - 4 5/8 11 1/8 = t

OpenStudy (anonymous):

thnks!

OpenStudy (anonymous):

np